# Topology of ultra-localized insulators and superconductors

### v1

@@ D $\alpha_{\text{Ch}_{1}}$ 00 00 00 00 00 00 $\alpha_{\text{Wi}_{9}}$ ${\text{Wi}_{9}}\mod 2$
Forgetting a symmetry — We start by proving the following statement: topologically non-trivial superconductors in class C (D) in dimensions $d=4n+1$ and $d=4n$, with $n$ an odd (even) integer, can be obtained by forgetting the time-reversal symmetry of a generating topological superconductor in class CII (BDI). Furthermore, topologically non-trivial insulators in class AII (AI) in dimensions $d=4n-1$, with $n$ odd (even) integer, can be obtained by forgetting the particle-hole symmetry of a generating topological superconductor in class DIII (CI). To prove these statements, it is sufficient to consider models for generators of the topological phases with the additional time-reversal symmetry or particle-hole symmetry and show that these remain generators of the topological phases if this additional symmetry constraint is lifted.
We start from a $(4n+1)$-dimensional topologically non-trivial superconductor in class CII (BDI) for $n$ odd (even), with the winding number ${\text{Wi}}=1$. The Hamiltonian of such system can be expressed using $24n+$ 2 (anticommuting) Dirac matrices,
@@
Next, we consider the generator of a $(4n-1)$-dimensional topological superconducting phase in class DIII (CI) for $n$ odd (even), which has winding number ${\text{Wi}}=1$. This generator can be expressed using the Dirac matrices
@@
which describe a $4n$-dimensional topological insulator in class AII (AI) with (generalized) Chern number $\text{Ch}=1$ for $n$ odd (even). Its first descendant is the $(4n-1)$-dimensional topological insulator in the same class
@@
We start from a $(4n+2)$-dimensional topologically non-trivial superconductor in class C (D) for $n$ odd (even), with Chern number $\text{Ch}=1$. The Hamiltonian of such system can be expressed using $34n+$ 3 Dirac matrices,
@@
Finally, the remaining topological superconductors in $d=3,7$ all have the winding number as their topological invariant. It is straightforward to see that they can be obtained by doubling a corresponding phase from class AIII. This way, we obtain the localizability of topological superconductors with even winding numbers in classes DIII and CI. For odd topological invariants in class DIII (CI), we note that forgetting particle-hole symmetry gives nontrivial topological insulators in class AII (AI), which are not localizable. Hence, $(4n-1)$-dimensional topological superconductors in class DIII (CI), with $n$ odd (even), are non-localizable if their winding number is an odd integer. For these classes, the group homomorphism $f$ is uniquely determined by requiring that its image consists of even winding numbers only, see Table 5.

### v2

@@ D $\alpha_{\text{Ch}_{1}}$ 0 0 0 0 0 0 $\alpha_{\text{Wi}_{9}}$ ${\text{Wi}_{9}}\mod 2$
Forgetting a symmetry — We start by proving the following statement: topologically non-trivial superconductors in class C (D) in dimensions $d=4n+1$ and $d=4n$, with $n$ an odd (even) integer, can be obtained by forgetting the time-reversal symmetry of a generating topological superconductor in class CII (BDI). Furthermore, topologically non-trivial insulators in class AII (AI) in dimensions d=4n1d=4n-1, with $n$ odd (even) integer, can be obtained by forgetting the particle-hole symmetry of a generating topological superconductor in class DIII (CI). To prove these statements, it is sufficient to consider models for generators of the topological phases with the additional time-reversal symmetry or particle-hole symmetry and show that these remain generators of the topological phases if this additional symmetry constraint is lifted.
We start from a $(4n+1)$-dimensional topologically non-trivial superconductor in class CII (BDI) for $n$ odd (even), with the winding number ${\text{Wi}}=1$. The Hamiltonian of such system can be expressed using $4n+2$ (anticommuting) Dirac matrices,
@@
Next, we consider the generator of a (4n1)(4n-1)-dimensional topological superconducting phase in class DIII (CI) for $n$ odd (even), which has winding number ${\text{Wi}}=1$. This generator can be expressed using the Dirac matrices
@@
which describe a $4n$-dimensional topological insulator in class AII (AI) with (generalized) Chern number $\text{Ch}=1$ for $n$ odd (even). Its first descendant is the (4n1)(4n-1)-dimensional topological insulator in the same class
@@
We start from a $(4n+2)$-dimensional topologically non-trivial superconductor in class C (D) for $n$ odd (even), with Chern number $\text{Ch}=1$. The Hamiltonian of such system can be expressed using $4n+3$ Dirac matrices,
@@
Finally, the remaining topological superconductors in $d=3,7$ all have the winding number as their topological invariant. It is straightforward to see that they can be obtained by doubling a corresponding phase from class AIII. This way, we obtain the localizability of topological superconductors with even winding numbers in classes DIII and CI. For odd topological invariants in class DIII (CI), we note that forgetting particle-hole symmetry gives nontrivial topological insulators in class AII (AI), which are not localizable. Hence, (4n1)(4n-1)-dimensional topological superconductors in class DIII (CI), with $n$ odd (even), are non-localizable if their winding number is an odd integer. For these classes, the group homomorphism $f$ is uniquely determined by requiring that its image consists of even winding numbers only, see Table 5.